# Berezin transform of Toeplitz operators on Bergman space with Békollé and Bonami weights

### v1

@@ 2.3. Duality
$\lesssim$ $\sum_{k=1}^{\infty}u(\Delta(z_{k},r))^{\frac{q(p+q^{\prime})}{pq^{\prime}}-1}\int_{\Delta(z_{k},r)}\left|f(\xi)K(z,\xi)\right|^{q}u(\xi)dA(\xi).$
By Fubini’s theorem, Lemma 2.6 and Lemma 2.1, we obtain
$\|T_{\mu}f\|_{A_{u}^{q}}^{q}$ $\lesssim$ $\int_{\mathbb{D}}\sum_{k=1}^{\infty}\int_{\Delta(z_{k},r)}u(\Delta(\xi,r))^{\frac{q(p+q^{\prime})}{pq^{\prime}}-1}\left|f(\xi)K(z,\xi)\right|^{q}u(\xi)dA(\xi)u(z)dA(z)$
$\lesssim$ $\int_{\mathbb{D}}\int_{\mathbb{D}}u(\Delta(\xi,r))^{\frac{q(p+q^{\prime})}{pq^{\prime}}-1}\left|f(\xi)K(z,\xi)\right|^{q}u(\xi)dA(\xi)u(z)dA(z)$
$\lesssim$ $u(\Delta(\xi,r))^{\frac{q(p+q^{\prime})}{pq^{\prime}}-1}\int_{\mathbb{D}}\left|f(\xi)\right|^{q}u(\xi)dA(\xi)\int_{\mathbb{D}}\left|K(z,\xi)\right|^{q}u(z)dA(z)$
$\lesssim$ $u(\Delta(\xi,r))^{\frac{q(p+q^{\prime})}{pq^{\prime}}-1}\|K_{\xi}\|_{A_{u}^{q}}^{q}\int_{\mathbb{D}}|f(\xi)|^{q}u(\xi)dA(\xi)$
$\simeq$ $u(\Delta(\xi,r))^{\frac{q(p+q^{\prime})}{pq^{\prime}}}\frac{1}{(1-|\xi|)^{2q}}\int_{\mathbb{D}}|f(\xi)|^{p}|f(\xi)|^{q-p}u(\xi)dA(\xi)$
$=$ $u(\Delta(\xi,r))^{\frac{q(p+q^{\prime})}{pq^{\prime}}}\frac{1}{(1-|\xi|)^{2q}}\int_{\mathbb{D}}|f(\xi)|^{p}|f(\xi)|^{p\cdot\frac{q-p}{p}}u(\xi)dA(\xi)$
$\lesssim$ $u(\Delta(\xi,r))^{\frac{q(p+q^{\prime})}{pq^{\prime}}}\frac{1}{(1-|\xi|)^{2q}}\int_{\mathbb{D}}|f(\xi)|^{p}\left(u(\Delta(\xi,r))^{-1}\int_{\mathbb{D}}|f(\xi)|^{p}u(\xi)dA(\xi)\right)^{\frac{q-p}{p}}u(\xi)dA(\xi)$
$\lesssim$ $u(\Delta(\xi,r))^{\frac{q(p+q^{\prime})}{pq^{\prime}}-\frac{q-p}{p}}\frac{1}{(1-|\xi|)^{2q}}\|f\|_{A_{u}^{p}}^{q-p}\int_{\mathbb{D}}|f(\xi)|^{p}u(\xi)dA(\xi)$
$=$ $\frac{u(\Delta(\xi,r))^{q}}{(1-|\xi|)^{2q}}\|f\|_{A_{u}^{p}}^{q}\simeq\|K_{\xi}\|_{A_{u}^{1}}^{q}\|f\|_{A_{u}^{p}}^{q},$
@@
4. case $1<q<p<\infty$
4.1. Proof of Theorem 1.3
@@
Since $1<q<p$, we have $\frac{p-q}{p}<1<1+q(1-\frac{1}{2p})$,
@@
5. essential norms
Let $\mathcal{K}$ be the set of all compact operators on a Banach space $\mathcal{B}$. For any bounded linear operator $T:\mathcal{B}\to\mathcal{B}$, the essential norm of $T$ is defined by
@@
5.1. case $0<p\leq q<\infty$
Theorem 5.1.
$\|T_{\mu}\|_{e}\simeq\limsup_{|z|\to 1}\widetilde{\mu_{t}}(z)u(\Delta(z,r))^{\frac{q-p}{pq}}\simeq\limsup_{|z|\to 1}\widehat{\mu_{r}}(z)u(\Delta(z,r))^{\frac{q-p}{pq}}$
@@
By Lemma 2.10, it suffices to prove that,
@@
5.2. case $0<q<p<\infty$
In this case, by Theorem 1.3, $T_{\mu}$ is bounded if and only if $T_{\mu}$ is compact. Consequently, the essential norm of $T_{\mu}$ is 0.
6. Schatten class
@@
Theorem 6.1.
$\int_{\mathbb{C}^{n}}h(C\widetilde{\mu}_{2}(z))\frac{u(\Delta(z,r))}{(1-|z|)^{4}}u(z)dA(z)<\infty$

### v2

@@ 2.3. Duality
$\lesssim$ $\sum_{k=1}^{\infty}u(\Delta(z_{k},r))^{\frac{q(p+q^{\prime})}{pq^{\prime}}-1}\int_{\Delta(z_{k},r)}\left|f(\xi)\right|^{q}\left|K(z,\xi)\right|^{q}u(\xi)dA(\xi)$
$\simeq$ $\sum_{k=1}^{\infty}\int_{\Delta(z_{k},r)}u(\Delta(\xi,r))^{\frac{q(p+q^{\prime})}{pq^{\prime}}-1}\left|f(\xi)\right|^{q}\left|K(z,\xi)\right|^{q}u(\xi)dA(\xi)$
By Fubini’s theorem, Lemma 2.6 and Lemma 2.1, we have
$\|T_{\mu}f\|_{A_{u}^{q}}^{q}$ $\lesssim$ $\int_{\mathbb{D}}\sum_{k=1}^{\infty}\int_{\Delta(z_{k},r)}u(\Delta(\xi,r))^{\frac{q(p+q^{\prime})}{pq^{\prime}}-1}\left|f(\xi)\right|^{q}\left|K(z,\xi)\right|^{q}u(\xi)dA(\xi)u(z)dA(z)$
$\lesssim$ $\int_{\mathbb{D}}\int_{\mathbb{D}}u(\Delta(\xi,r))^{\frac{q(p+q^{\prime})}{pq^{\prime}}-1}\left|f(\xi)\right|^{q}\left|K(z,\xi)\right|^{q}u(\xi)dA(\xi)u(z)dA(z)$
$=$ $\int_{\mathbb{D}}u(\Delta(\xi,r))^{\frac{q(p+q^{\prime})}{pq^{\prime}}-1}\|K_{\xi}\|_{A_{u}^{q}}^{q}\left|f(\xi)\right|^{q}u(\xi)dA(\xi)$
$=$ $\int_{\mathbb{D}}u(\Delta(\xi,r))^{\frac{q(p+q^{\prime})}{pq^{\prime}}-1}\|K_{\xi}\|_{A_{u}^{q}}^{q}|f(\xi)|^{p}|f(\xi)|^{p\cdot\frac{q-p}{p}}u(\xi)dA(\xi)$
$\lesssim$ $\int_{\mathbb{D}}u(\Delta(\xi,r))^{\frac{q(p+q^{\prime})}{pq^{\prime}}-1}\|K_{\xi}\|_{A_{u}^{q}}^{q}|f(\xi)|^{p}$
$\cdot\left(u(\Delta(\xi,r))^{-1}\int_{u(\Delta(\xi,r))}|f(w)|^{p}u(w)dA(w)\right)^{\frac{q-p}{p}}u(\xi)dA(\xi)$
$\lesssim$ $\|f\|_{A_{u}^{p}}^{q-p}\int_{\mathbb{D}}u(\Delta(\xi,r))^{q-1}\|K_{\xi}\|_{A_{u}^{q}}^{q}|f(\xi)|^{p}u(\xi)dA(\xi)$
$\simeq$ $\|f\|_{A_{u}^{p}}^{q-p}\int_{\mathbb{D}}\frac{u(\Delta(\xi,r))^{q}}{(1-|\xi|)^{2q}}|f(\xi)|^{p}u(\xi)dA(\xi)$
$\simeq$ $\|f\|_{A_{u}^{p}}^{q-p}\int_{\mathbb{D}}\|K_{\xi}\|_{A_{u}^{1}}^{q}|f(\xi)|^{p}u(\xi)dA(\xi)$
$\leq$ $\sup_{\xi\in\mathbb{D}}\|K_{\xi}\|_{A_{u}^{1}}^{q}\|f\|_{A_{u}^{p}}^{q}$
@@
5. Proof of Theorem 1.3
@@
When $q<p$, we have $0<\frac{p-q}{p}<1+q(1-\frac{1}{2p})$,
@@
6. Proof of Theorem 1.4 and Theorem 1.5
First, we recall some definitions about essential norm. Let $\mathcal{K}$ be the set of all compact operators on a Banach space $\mathcal{B}$. For any bounded linear operator $T:\mathcal{B}\to\mathcal{B}$, the essential norm of $T$ is defined by
@@
Now we give the proof of Theorem 1.4.
@@
When $0<p\leq q<\infty$, by Lemma 2.10, it suffices to prove that,
@@
And we note that when $0<q<p<\infty$, by Theorem 1.3, $T_{\mu}$ is bounded if and only if $T_{\mu}$ is compact. Consequently, the essential norm of $T_{\mu}$ is 0.
@@
Then, we give the proof of Theorem 1.5.